# Long Chain-of-Thought for Reasoning Large Language Models

### v3

@@ Abstract
Recent advancements in reasoning with large language models (RLLMs), such as OpenAI-O1 and DeepSeek-R1, have demonstrated their impressive capabilities in complex domains like mathematics and coding. A central factor in their success lies in the application of long chain-of-thought (Long CoT) characteristics, which enhance reasoning abilities and enable the solution of intricate problems. However, despite these developments, a comprehensive survey on Long CoT is still lacking, limiting our understanding of its distinctions from traditional short chain-of-thought (Short CoT) and complicating ongoing debates on issues like "overthinking" and "test-time scaling." This survey seeks to fill this gap by offering a unified perspective on Long CoT. (1) We first distinguish Long CoT from Short CoT and introduce a novel taxonomy to categorize current reasoning paradigms. (2) Next, we explore the key characteristics of Long CoT: deep reasoning, extensive exploration, and feasible reflection, which enable models to handle more complex tasks and produce more efficient, coherent outcomes compared to the shallower Short CoT. (3) We then investigate key phenomena such as the emergence of Long CoT with these characteristics, including overthinking, and test-time scaling, offering insights into how these processes manifest in practice. (4) Finally, we identify significant research gaps and highlight promising future directions, including the integration of multi-modal reasoning, efficiency improvements, and enhanced knowledge frameworks. By providing a structured overview, this survey aims to inspire future research and further the development of logical reasoning in artificial intelligence [^1] Our logo refers to a cute cartoon image - Snake Puppy. Header Image is inspired by Yaoting et al. [759].
@@
1 Introduction
In recent years, the emergence of reasoning large language models (RLLMs) such as OpenAI O1 [240] and DeepSeek R1 [179] has sparked a growing body of research into Long Chain-of-Thought (Long CoT) reasoning, greatly improving their mathematical reasoning, programming tasks, and multidisciplinary knowledge reasoning capabilities [549, 777, 572, 60, 68, 761, 155, 880], even passing Turing Test [263], as shown in Figure 1. This shift marks a significant departure from traditional approaches to task handling in large language models (LLMs) [905, 492, 494, 476]. Unlike the shorter chain-of-thought (Short CoT) used in traditional LLMs, Long CoT reasoning entails a more detailed, iterative process of exploration and reflection within a given problem space by test-time scaling [338, 584, 413]. This process has led to notable advancements in mathematical and logical reasoning, as well as in exploring how supervised fine-tuning (SFT) and reinforcement learning (RL) techniques can enhance the learning and exploration of extended reasoning chains [495, 435].
However, there is no comprehensive survey to systematically understand the main factors and recent efforts of Long CoT for RLLMs, which hinders the development of RLLMs. As a result, there are ongoing debates about the effectiveness of simple "test-time scaling" for Longer CoT [686, 386] versus the argument that "over-thinking" from excessively long scaling can harm LLMs and introduce unnecessary complexity [87, 114, 285]. Moreover, some researchers argue that, when solving specific problems, there is no clear relationship between length and accuracy [701].
To address this gap, we provide an extensive and comprehensive survey of Long CoT. Specifically, as illustrated in Figure 2, we first define and examine the distinctions between Long CoT and traditional Short CoT, focusing on the following key aspects: (1) Deep Reasoning, which requires a sufficient depth of logical processing to manage an extensive set of reasoning nodes; (2) Extensive Exploration, which involves generating parallel uncertain nodes and transitioning from known to unknown logic; and (3) Feasible Reflection, which involves feedback and refinement of logical connections. These characteristics enable Long CoT paradigms to integrate more intricate reasoning and accommodate a broader range of logical structures, ultimately leading to more efficient and coherent outcomes. Subsequently, we systematically explore the underlying explanations for key phenomena associated with Long CoT, such as its emergence, the overthinking phenomenon, inference time scaling during testing, and the "Aha Moment," among others. To our knowledge, This is the first comprehensive survey dedicated to these specific topics. Finally, considering the extensive body of literature, we highlight promising areas for future research and suggest valuable open-resource frameworks and datasets that can serve as a foundation for future investigations.
@@ 1 Introduction
  • Explanation of Hot Phenomena: We systematically investigate the notable phenomena associated with Long CoT reasoning, such as overthinking, inference test-time scaling, and the “Aha Moment”, offering valuable insights into the cognitive processes involved in complex reasoning.
@@ 2.1 Overview of Short CoT
where $n_{1}$ to $n_{k}$ represent a sequence of logical nodes, which naturally satisfy that $\forall i,n_{i}\rightarrow n_{i+1}$. Here, $\mathcal{B}_{s}$ denotes the upper boundary on the number of reasoning nodes, as defined by Chen et al. [78]. In this paradigm, the reasoning progresses sequentially from one node to the next, with minimal revisitation of previous nodes and little exploration of alternative logical paths.
@@ 2.2 Overview of Long CoT
In contrast, Long CoT involves deeper reasoning, reflective analysis, and a broader exploration of logical structures. It facilitates reasoning across a wider range of logical steps, addressing both known and unknown elements of a problem [150]. Formally, Long CoT expands the constraints outlined in Equation 1 based on either an explicit or implicit tree structure. As shown in Figure 3, we will now discuss these key differences in detail.
@@ 2.2.1 Deep Reasoning for Long CoT
where $\mathcal{B}_{l}$ represents the upper boundary for Long CoT reasoning, which can accommodate much more intricate reasoning nodes compared to the smaller boundary $\mathcal{B}_{s}$ for Short CoT. The larger boundary $\mathcal{B}_{l}$ alleviates issues related to insufficient depth in reasoning, thereby reducing the risk of generating unresolved answers or hallucinated responses in short-form reasoning.
@@ 2.2.2 Extensive Exploration for Long CoT
$j=1\Leftrightarrow\forall i\leq k,n_{i}\rightarrow n_{i+j}\rightarrowtail\exists m,\forall i,j\leq m,n_{i}\rightarrow n_{i+j},$ (3)
@@ 2.2.3 Feasible Reflection for Long CoT
As shown by Figure 2, Long CoT involves revisiting previous logical nodes to verify their connections are valid and accurate, and then correcting them or selecting an alternative logical path. Formally, feasible reflection relaxes the third constraint in Equation 1, expressed as follows:
@@
where this condition indicates that, for a logical node $n_{j-1}$, the subsequent node is not limited to the original next node $\hat{n}_{j}$. Instead, it may transition to $n_{i}$ (i.e., the next reasoning node becomes $n_{j}$, where $n_{j}=n_{i}$). Practically, reflection implementation consists of two components:
@@
Feedback refers to evaluating both overall and intermediate outputs for correctness and quality, also known as critique or verification. It can be derived from external sources, validation checks, or by reflecting on prior conclusions within the reasoning process. Formally, at each step $n_{i}$, a verification process $\mathcal{V}_{i}$ ensures the correctness, feasibility, and consistency of the reasoning. If an issue is identified, the process redirects $n_{i}$ to the nearest correct node $n_{j}$, where $j<i$. This relationship is formalized as:
@@
where $\texttt{CoT}_{L}^{i}=\{n_{1},\dots,n_{i}\}$ represents the current logic path up to the $i$-th logic node for Long CoT.
@@
This involves adjusting intermediate steps or modifying the logical flow to correct inconsistencies or address gaps based on the given feedback. This process can be expressed mathematically as follows:
@@
where $\widetilde{n}_{i+1}$ represents the refined version of the subsequent reasoning node $n_{i+1}$, according to the current logic $\texttt{CoT}_{L}^{i}$, feedback result $\mathcal{F}_{i}$, and previous reasoning node $n_{j}$.
@@
2.2.4 Unified Discussion
The Long CoT discussed here represents a unified reasoning system that integrates the three key characteristics outlined earlier. In contrast, during the Short CoT era, these abilities developed independently. As shown in Figure 2, early efforts focus on deep reasoning within traditional CoT paradigms, followed by the gradual introduction of reflective mechanisms based on human-designed pipeline. Exploration capabilities are then added, and these three components are merged to form the modern concept of Long CoT, which aims for a unified enhancement of reasoning.
@@ 2.2.4 Unified Discussion
Key Difference: Unified Capabilities It is important to highlight that Long CoT integrates the following three distinct capabilities to perform complex reasoning. In contrast, traditional Short CoT optimization typically focuses on only one of these characteristics.
@@ 3.1.1 Long CoT External Behavior Analysis
Research shows that contextual examples improve large models’ generative abilities by guiding the formation of reasoning chains [804, 532, 336, 272, 418]. Wang et al. [609] and Lippmann and Yang [367] demonstrate that these examples standardize reasoning chain generation relevant to the answers both in in-context-learning and supervised-finetuning. In an experiment by Madaan et al. [424], removing problem-specific entities from contextual examples, while retaining only the logical structure, led to similar performance as using complete examples, highlighting the logical structure imitation of Long CoT during inference.
More recently, Stechly et al. [544] and Wang and Zhou [651] have shown that modifying the decoding process or designing specific prompts can activate the Long CoT within pre-trained models. They propose that CoT is embedded during pre-training and requires specific activation [745]. Further, Sadr et al. [512] focus the Long CoT source from the training data, and build on this with the notion of “model attribution”, to specifically identify the training data most influential for specific outputs. Building on this, Guo et al. [179] and Xie et al. [701] investigate using rule-based reinforcement learning to directly activate Long CoT during pre-training, aiming to enhance performance [697]. Furthermore, Gandhi et al. [150] identify four key cognitive behaviors, including verification, backtracking, sub-target setting, and backlinking, which successfully facilitate Long CoT. Qwen [733] inherently demonstrates these behaviors, which can be easily triggered by rule-based reinforcement. In contrast, Llama [134] lacks these capabilities and thus requires example-based reinforcement learning to improve significantly. Moreover, Wang et al. [648] identify a pretraining scaling law that explains how increasing calculation size in RLLMs enhances their reasoning capabilities.
Figure 4: Analysis of the six classic phenomena of Long CoT external behavior: (a) emergence of Long CoT in current RLLMs; (b) reasoning boundaries and limitations of current Long CoT systems; (c) overthinking caused by scaling beyond RLLMs’ reasoning boundaries, leading to performance decay; (d) test-time scaling, discussing mainstream scaling methods, corresponding scaling laws and their limitations; (e) use of process reward model (PRM) or outcome reward model (ORM); (f) exploration of the "aha" moment and its underlying causes.
@@
Recent research has highlighted the upper bounds and limitations of RLLMs across various reasoning tasks [236, 220, 541, 203, 143]. Specifically, Bi et al. [46] investigate these bounds in code generation, showing that RLLMs struggle with tasks that exceed certain complexity thresholds, especially when imitating Long CoT samples of varying complexity. In the context of upper-bound performance, Merrill and Sabharwal [433] and Li et al. [345] focus on single-step arithmetic tasks, concluding that model performance is constrained by input length. Moreover, Feng et al. [139] proposes a mathematical model indicating that fixed-size models cannot produce accurate numerical answers beyond specific limits. However, increasing the number of reasoning steps improves a model’s ability to solve more complex problems.
Inspired by these explorations, Chen et al. [78] first define the “reasoning boundary” phenomenon and quantify these limits, showing that surpassing an RLLM’s reasoning capacity leads to performance decline. Similarly, Zhou et al. [895] introduce GSM-Infinite, linking different upper limits to accuracy levels. Chen et al. [78] also examine the interaction between these boundaries across tasks of varying complexity, providing insights into the effectiveness of Long CoT strategies [861]. Moreover, Amiri et al. [10] propose a “tight lower bound” for Long CoT further guiding reasoning error reductions. Further, Baeumel et al. [25] suggest that due to its reliance on a single-digit lookahead heuristic, there are inherent boundaries in performing addition with multiple operands, which thus hinders the fundamental limitation of LLMs in scaling to more complex numerical reasoning.
@@
Research has highlighted the overthinking phenomenon [87, 260, 454, 114, 285], where performance improves with longer reasoning chains up to a threshold, after which it declines. In contrast, Xie et al. [701] and Ma et al. [420] find no significant correlation between reasoning length and accuracy. To explain this, one line of research suggests that Long CoT strategies, like avoiding “snowball errors” [148]. Alternatively, Chen et al. [78], Wolf et al. [677] highlight a performance drop when the reasoning boundaries are exceeded, providing an explanation for the overthinking phenomenon. This suggests that reasoning length and logical complexity should be kept below a certain boundary [857]. Building on this, Wu et al. [687] mathematically determine the feasible reasoning length for Long CoT. Finally, Chen et al. [80] introduces Ohm’s law of Long CoT, which accurately predicts and controls performance.
Inference Test-Time Scaling Phenomenon
@@
Recent advances in inference-time scaling algorithms [413, 672] have garnered significant attention, particularly for their ability to extend reasoning length and improve performance [413]. Specifically, Brown et al. [50] identify a phenomenon called "Large Language Monkeys," in which a series of reasoning tasks show that with enough trials, a correct result can be achieved. Additionally, o1 [240] and R1 [179] demonstrated that directly scaling the length of model inference improves final performance.
@@
To understand inference test-time scaling, we will discuss these two paradigms: (1) Vertical Scaling: Vertical scaling involves increasing the reasoning path length. While this can enhance performance, studies by Jin et al. [260] show that, beyond a certain point, longer reasoning paths can degrade performance due to error accumulation. They suggest an optimal path length that depends on the model’s capabilities and task complexity [13, 520, 28]. Furthermore, Chen et al. [78] and Wu et al. [687] explain that excessive exploration lengths beyond the RLLM’s inherent reasoning boundary lead to performance decay, which guides RLLMs for deeper reasoning capabilities [29]. (2) Parallel Scaling: Parallel scaling involves performing multiple reasoning steps and verifying the results. While it shows promise, Parashar et al. [462] and Wang et al. [656] argue that simply increasing inference time does not guarantee improved performance. Wu et al. [686] show that the computational FLOPs $N$ of inference are correlated with the lower bound of performance error, which scales with $\log N$. Additionally, Chen et al. [80] establish an upper bound for parallel scaling, showing that RLLMs cannot exceed Pass@k verification through various verifiers. They further argue that sampling optimization cannot exceed the model’s internal reasoning limitations, demonstrating that for $N$ samples, accuracy is proportional to $\frac{m}{(k/\log N+b)^{2}}$, where $m$, $n$, and $b$ are model-dependent constants.
PRM v.s. ORM Phenomenon
@@
As RLLMs evolve, it is crucial to differentiate between process supervision and outcome supervision, two key reinforcement learning approaches for complex reasoning tasks, which is essential [712, 145, 841]. While process supervision is intuitively advantageous for long-term reward assignments, the exact relationship between the two remains unclear. It is commonly believed that process supervision is more challenging due to the trajectory-level coverage problem, which demands significant effort to collect fine-grained supervision data [871, 538]. Additionally, PRM faces the issue of reward hacking [11, 121, 453, 27, 322], where agents exploit flaws in the reward function to produce unintended behaviors [179]. Addressing this to surpass rule-based reward systems has become an important research area [179, 701, 472]. Furthermore, Lampinen et al. [295] and Tan [559] establish a causal link between intermediate steps and final answers in qualitative experiments. Building on this, Jia et al. [249] demonstrate that, under the standard data coverage assumption, reinforcement learning with outcome supervision is not statistically more challenging than process supervision, aside from polynomial factors.
@@
Earlier, Guo et al. [179] demonstrated that direct RL using rule-based rewards can trigger the aha moment, fostering natural self-reflection without supervision. Following this, Team [571], Xie et al. [701] replicated this phenomenon. Further, Zhou et al. [886] and Meng et al. [432] further extend this phenomenon to multimodal scenarios. However, Liu et al. [392] argue that the aha moment may not emerge in R1-Zero-like training. Instead, they observe that self-reflection patterns, such as superficial self-reflection (SSR), appear at epoch 0, the stage of base models. In this case, self-reflections do not necessarily lead to correct answers. Upon closer examination of R1-Zero training via RL, they find that the increasing response length results not from self-reflection, but from RL optimizing well-designed rule-based rewards. Moreover, Yang et al. [744] demonstrate that the "aha moment" is externally marked by increased use of anthropomorphic language during self-reflection and a dynamic adjustment of uncertainty in response to problem difficulty. This process enables the model to maintain reasoning without succumbing to "Reasoning Collapse." Internally, it is characterized by a clear distinction between anthropomorphic traits and logical reasoning, with anthropomorphic language intensifying as the problem becomes more complex.

### v4

@@ Abstract
Recent advancements in reasoning with large language models (RLLMs), such as OpenAI-o1 and DeepSeek-R1, have demonstrated their impressive capabilities in complex domains like mathematics and coding. A central factor in their success lies in the application of long chain-of-thought (Long CoT) characteristics, which enhance reasoning abilities and enable the solution of intricate problems. However, despite these developments, a comprehensive survey on Long CoT is still lacking, limiting our understanding of its distinctions from traditional short chain-of-thought (Short CoT) and complicating ongoing debates on issues like “overthinking” and “inference-time scaling”. This survey seeks to fill this gap by offering a unified perspective on Long CoT. Specifically, (1) We first distinguish Long CoT from Short CoT and introduce a novel taxonomy to categorize current reasoning paradigms. (2) Next, we explore the key characteristics of Long CoT: deep reasoning, extensive exploration, and feasible reflection, which enable models to handle more complex tasks and produce more efficient, coherent outcomes compared to the shallower Short CoT. (3) We then investigate key phenomena such as the emergence of Long CoT with these characteristics, including overthinking, and inference-time scaling, offering insights into how these processes manifest in practice. (4) Finally, we identify significant research gaps and highlight promising future directions, including the integration of multi-modal reasoning, efficiency improvements, and enhanced knowledge frameworks. By providing a structured overview, this survey aims to inspire future research and further the development of reasoning large language models  [^1] Our logo refers to a cute cartoon image - Snake Puppy. Header Image is inspired by Yaoting et al. [923].
1 Introduction
In recent years, as shown in Figure 1, the emergence of reasoning large language models (RLLMs) such as OpenAI o1 [289] and DeepSeek R1 [213] has sparked a growing body of research into Long Chain-of-Thought (Long CoT) reasoning, greatly improving their mathematical reasoning, programming tasks, and multidisciplinary knowledge reasoning capabilities [668, 943, 694, 75, 925, 187, 1071], even passing Turing Test [315]. This shift marks a significant departure from traditional approaches to task handling in large language models (LLMs) [1104, 593, 596, 573]. Unlike the shorter chain-of-thought (Short CoT) used in traditional LLMs, Long CoT reasoning entails a more detailed, iterative process of exploration and reflection within a given problem space by inference-time scaling [398, 706, 501]. This process has led to notable advancements in mathematical and logical reasoning, as well as in exploring how supervised fine-tuning (SFT) and reinforcement learning (RL) techniques can enhance the learning and exploration of extended reasoning chains [597, 526].
@@ 1 Introduction
However, there is no comprehensive survey to systematically understand the main factors and recent efforts of Long CoT for RLLMs, which hinders the development of RLLMs. As a result, there are ongoing debates about the effectiveness of simple “inference-time scaling” for Longer CoT [831, 463] versus the argument that “over-thinking” from excessively long scaling can harm LLMs and introduce unnecessary complexity [98, 136, 338]. Moreover, some researchers argue that, when solving specific problems, there is no clear relationship between length and accuracy [852].
To address this gap, we provide an extensive and comprehensive survey of Long CoT. Specifically, as illustrated in Figure 2, we first define and examine the distinctions between Long CoT and traditional Short CoT, focusing on the following key aspects: (1) Deep Reasoning, which requires a sufficient depth of logical processing to manage an extensive set of logical nodes; (2) Extensive Exploration, which involves generating parallel uncertain nodes and transitioning from known to unknown logic; and (3) Feasible Reflection, which involves feedback and refinement of logical connections. These characteristics enable Long CoT paradigms to integrate more intricate reasoning and accommodate a broader range of logical structures, ultimately leading to more efficient and coherent outcomes. Subsequently, we systematically explore the underlying explanations for key phenomena associated with Long CoT, such as its emergence, the overthinking phenomenon, inference-time scaling during testing, and the "Aha Moment," among others. To our knowledge, This is the first comprehensive survey dedicated to these specific topics. Finally, considering the extensive body of literature, we highlight promising areas for future research and suggest valuable open-resource frameworks and datasets that can serve as a foundation for future investigations.
@@
  • Explanation of Hot Phenomena: We systematically investigate the notable phenomena associated with Long CoT reasoning, such as overthinking, inference-time scaling, and the “Aha Moment”, offering valuable insights into the cognitive processes involved in complex reasoning.
@@ 2.1 Overview of Short CoT
where $n_{1}$ to $n_{k}$ represent a sequence of logical nodes, which naturally satisfy that $\forall i,n_{i}\rightarrow n_{i+1}$. Here, $\mathcal{B}_{s}$ denotes the upper boundary on the number of logical nodes, as defined by Chen et al. [85]. In this paradigm, the reasoning progresses sequentially from one node to the next, with minimal revisitation of previous nodes and little exploration of alternative logical paths.
@@ 2.2 Overview of Long CoT
In contrast, Long CoT involves deeper reasoning, reflective analysis, and a broader exploration of logical structures. It facilitates reasoning across a wider range of logical steps, addressing both known and unknown elements of a problem [181, 825]. Building on this, Long CoT expands upon the constraints presented in Equation 1 based on tree structures by incorporating three critical components: deep reasoning, exploration, and reflection.
These components play distinct yet complementary roles in enhancing cognitive processes. Deep reasoning ensures each logical step is executed rigorously, even within complex structures, fostering robust logic across intricate relationships. Exploration encourages the identification of new pathways, revealing potential avenues that may not be immediately obvious. Reflection enables iterative analysis and reassessment of conclusions, allowing reasoning to evolve throughout problem-solving. By distinguishing these three categories, Long CoT enhances its ability to address a broader range of problems with precision and depth. As shown in Figure 3, we will now discuss these key differences in detail.
@@ 2.2.1 Deep Reasoning for Long CoT
where $\mathcal{B}_{l}$ represents the upper boundary for Long CoT reasoning, which can accommodate much more intricate logical nodes compared to the smaller boundary $\mathcal{B}_{s}$ for Short CoT. The larger boundary $\mathcal{B}_{l}$ alleviates issues related to insufficient depth in reasoning, thereby reducing the risk of generating unresolved answers or hallucinated responses in short-form reasoning.
@@ 2.2.2 Extensive Exploration for Long CoT
$j=1\Leftrightarrow\forall i\leq k,n_{i}\rightarrow n_{i+j}\rightarrowtail\exists m,\forall i,\forall j\leq m,n_{i}\rightarrow n_{i+j},$ (3)
@@ 2.2.3 Feasible Reflection for Long CoT
As shown by Figure 2, Long CoT involves revisiting previous logical nodes to verify their connections are valid and accurate, and then correcting them or selecting an alternative logical path. Formally, feasible reflection relaxes the third constraint in Equation 1, which originally requires acyclic reasoning such that $n_{i}\neq n_{j}$ for all $i\neq j\leq k$. In contrast, feasible reflection permits the reasoning path to return to a previously visited node, captured as:
@@
where this condition indicates that, for a logical node $n_{j-1}$, the subsequent node is not limited to the original next node $\hat{n}_{j}$. Instead, it may transition to $n_{i}$ (i.e., the next logical node becomes $n_{j}$, where $n_{j}=n_{i}$). Practically, reflection implementation consists of two components:
@@
refers to evaluating both overall and intermediate outputs for correctness and quality, also known as critique or verification. It can be derived from external sources, validation checks, or by reflecting on prior conclusions within the reasoning process. Formally, at each step $n_{i}$, a verification process $\mathcal{V}_{i}$ ensures the correctness, feasibility, and consistency of the reasoning. If an issue is identified, the process redirects $n_{i}$ to the nearest correct node $n_{j}$, where $j<i$. This relationship is formalized as:
@@
where $\texttt{CoT}_{L}^{i}=\{n_{1},\dots,n_{i}\}$ represents the current logical path up to the $i$-th logical node for Long CoT.
@@
involves adjusting intermediate steps or modifying the logical flow to correct inconsistencies or address gaps based on the given feedback. This process can be expressed mathematically as follows:
@@
where $\widetilde{n}_{i+1}$ represents the refined version of the subsequent logical node $n_{i+1}$, according to the current logic $\texttt{CoT}_{L}^{i}$, feedback result $\mathcal{F}_{i}$, and previous logical node $n_{j}$.
@@
2.2.4 Unified Application and Development History of Three Capabilities
The Long CoT discussed here represents a unified reasoning system that seamlessly integrates and applies three key capabilities: deep reasoning, reflective mechanisms, and exploration capabilities. In contrast, during the Short CoT era, these capabilities developed independently, each evolving in isolation.
As shown in Figure 2, early efforts primarily focused on enhancing deep reasoning within traditional CoT paradigms. This was followed by the gradual introduction of reflective mechanisms, which were initially based on human-designed pipelines. Over time, exploration capabilities were added, and these components were ultimately merged, giving rise to the modern concept of Long CoT, a unified approach to reasoning that seeks to enhance all three capabilities in harmony.
@@ 2.2.4 Unified Application and Development History of Three Capabilities
Key Difference: Unified Application of Three Capabilities It is important to highlight that Long CoT integrates these three distinct capabilities to perform complex reasoning. In contrast, traditional Short CoT optimization typically focuses on only one of these characteristics.
@@ 3.1.1 Long CoT External Behavior Analysis
Research shows that contextual examples improve large models’ generative abilities by guiding the formation of reasoning chains [974, 644, 396, 324, 509, 814, 979, 1098]. Wang et al. [732] and Lippmann and Yang [438] demonstrate that these examples standardize reasoning chain generation relevant to the answers both in in-context-learning and supervised-finetuning. In an experiment by Madaan et al. [515], removing problem-specific entities from contextual examples, while retaining only the logical structure, led to similar performance as using complete examples, highlighting the logical structure imitation of Long CoT during inference. From a learning perspective, Ye et al. [927] analyzes and reveals the three-stage developmental trajectory of Long CoT: early memorization, followed by in-distribution generalization, and ultimately cross-distribution generalization, thereby enabling the model to exhibit Long CoT capabilities.
More recently, Stechly et al. [660] and Wang and Zhou [785] have shown that modifying the decoding process or designing specific prompts can activate the Long CoT within pre-trained models. They propose that CoT is embedded during pre-training and requires specific activation [906]. Further, Sadr et al. [616] focus the Long CoT source from the training data, and build on this with the notion of “model attribution”, to specifically identify the training data most influential for specific outputs. Building on this, Guo et al. [213] and Xie et al. [852] investigate using rule-based reinforcement learning to directly activate Long CoT during pre-training, aiming to enhance performance [847]. Furthermore, Gandhi et al. [181] identify four key cognitive behaviors, including verification, backtracking, sub-target setting, and backward chaining, which successfully facilitate Long CoT. Qwen series models [891] inherently demonstrate these behaviors, which can be easily triggered by rule-based reinforcement. In contrast, the models of Llama series [158] lack these capabilities and thus requires example-based reinforcement learning to improve significantly [61]. Moreover, Wang et al. [782] identify a pretraining scaling law that explains how increasing calculation size in RLLMs enhances their reasoning capabilities. Wang et al. [766] further explore the scaling law of Long CoT, showing that more fine-grained Long CoT granularity leads to more efficient and effective generalization performance.
Figure 4: Analysis of the six classic phenomena of Long CoT external behavior: (a) emergence of Long CoT in current RLLMs; (b) reasoning boundaries and limitations of current Long CoT systems; (c) overthinking caused by scaling beyond RLLMs’ reasoning boundaries, leading to performance decay; (d) inference-time scaling, discussing mainstream scaling methods, corresponding scaling laws and their limitations; (e) use of process reward model (PRM) or outcome reward model (ORM); (f) exploration of the "aha" moment and its underlying causes.
@@
Recent research has highlighted the upper bounds and limitations of RLLMs across various reasoning tasks [285, 266, 657, 244, 173, 236]. Specifically, Bi et al. [49] investigate these bounds in code generation, showing that RLLMs struggle with tasks that exceed certain complexity thresholds [574], especially when imitating Long CoT samples of varying complexity. In the context of upper-bound performance, Merrill and Sabharwal [524] and Li et al. [409] focus on single-step arithmetic tasks, concluding that model performance is constrained by input length. Moreover, Feng et al. [166] proposes a mathematical model indicating that fixed-size models cannot produce accurate numerical answers beyond specific limits. However, increasing the number of reasoning steps improves a model’s capability requirements to solve more complex problems.
Inspired by these explorations, Chen et al. [85] first define the “reasoning boundary” phenomenon and quantify these limits, showing that surpassing an RLLM’s reasoning capacity leads to performance decline [87]. Similarly, Zhou et al. [1088] introduce GSM-Infinite, linking different upper limits to accuracy levels. Chen et al. [85] also examine the interaction between these boundaries across tasks of varying complexity, providing insights into the effectiveness of Long CoT strategies [1044]. Moreover, Amiri et al. [11] propose a “tight lower bound” for Long CoT further guiding reasoning error reductions. Further, Baeumel et al. [27] suggest that due to its reliance on a single-digit lookahead heuristic, there are inherent boundaries in performing addition with multiple operands, which thus hinders the fundamental limitation of LLMs in scaling to more complex numerical reasoning. Liu et al. [460] further investigate the role of reinforcement learning in expanding these reasoning boundaries instead of relying solely on pretraining capabilities.
@@
Research has highlighted the overthinking phenomenon [98, 311, 548, 136, 338, 569], where performance improves with longer reasoning chains up to a threshold, after which it declines. In contrast, Xie et al. [852] and Ma et al. [511] find no significant correlation between reasoning length and accuracy. To explain this, one line of research suggests that Long CoT strategies [20, 418], like avoiding “snowball errors” [179]. Alternatively, Chen et al. [85], Wolf et al. [819] highlight a performance drop when the reasoning boundaries are exceeded, providing an explanation for the overthinking phenomenon. This suggests that reasoning length and logical complexity should be kept below a certain boundary [1039]. Building on this, Wu et al. [834] mathematically determine the feasible reasoning length for Long CoT. Finally, Chen et al. [88] introduces Ohm’s law of Long CoT, which accurately predicts and controls performance.
Inference-Time Scaling Phenomenon
@@
Recent advances in inference-time scaling algorithms [501, 811] have garnered significant attention, particularly for their ability to extend reasoning length and improve performance [501, 432, 841]. Specifically, Brown et al. [53] identify a phenomenon called "Large Language Monkeys," in which a series of reasoning tasks show that with enough trials, a correct result can be achieved. Additionally, o1 [289] and R1 [213] demonstrated that directly scaling the length of model inference improves final performance.
@@
To understand inference-time scaling, we will discuss these two paradigms: (1) Sequential Scaling: Sequential scaling involves increasing the reasoning path length. While this can enhance performance, studies by Jin et al. [311] show that, beyond a certain point, longer reasoning paths can degrade performance due to error accumulation. They suggest an optimal path length that depends on the model’s capabilities and task complexity [14, 625, 30]. Furthermore, Chen et al. [85] and Wu et al. [834] explain that excessive exploration lengths beyond the RLLM’s inherent reasoning boundary lead to performance decay, which guides RLLMs for deeper reasoning capabilities [31]. (2) Parallel Scaling: Parallel scaling involves performing multiple reasoning steps and verifying the results. While it shows promise, Parashar et al. [557] and Wang et al. [790] argue that simply increasing inference time does not guarantee improved performance. Wu et al. [831] show that the computational FLOPs $N$ of inference are correlated with the lower bound of performance error, which scales with $\log N$. Additionally, Chen et al. [88] establish an upper bound for parallel scaling, showing that RLLMs cannot exceed Pass@k verification through various verifiers. They further argue that sampling optimization cannot exceed the model’s internal reasoning limitations, demonstrating that for $N$ samples, accuracy is proportional to $\frac{m}{(k/\log N+b)^{2}}$, where $m$, $n$, and $b$ are model-dependent constants.
PRM & ORM Selection Phenomenon
@@
As RLLMs evolve, it is crucial to navigate the debate around the selection between process supervision and outcome supervision, two key reinforcement learning paradigms for complex reasoning tasks. The phenomenon of choosing between these two approaches has become a pivotal issue, as it is essential to differentiate and decide which supervision strategy is more suitable for specific tasks [865, 175, 1019]. While process supervision is intuitively advantageous for long-term reward assignments, the exact relationship between the two approaches remains unclear. It is commonly believed that process supervision is more challenging due to the trajectory-level coverage problem, which demands significant effort to collect fine-grained supervision data [1060, 652]. Additionally, PRM faces the issue of reward hacking [12, 144, 547, 29, 379], where agents exploit flaws in the reward function to produce unintended behaviors [213]. Addressing this to surpass rule-based reward systems has become an important research area [213, 852, 568]. Furthermore, Lampinen et al. [349] and Tan [680] establish a causal link between intermediate steps and final answers in qualitative experiments. Building on this, Jia et al. [298] demonstrate that, under the standard data coverage assumption, reinforcement learning with outcome supervision is not statistically more challenging than process supervision, aside from polynomial factors.
@@
Earlier, Guo et al. [213] demonstrated that direct RL using rule-based rewards can trigger the aha moment, fostering natural self-reflection without supervision [162]. Following this, Team [693], Xie et al. [852] replicate this phenomenon. Further, Zhou et al. [1077] and Meng et al. [523] further extend this phenomenon to multimodal scenarios. However, Liu et al. [475] argue that the aha moment may not emerge in R1-Zero-like training. Instead, they observe that self-reflection patterns, such as superficial self-reflection (SSR), appear at epoch 0, the stage of base models. In this case, self-reflections do not necessarily lead to correct answers. Upon closer examination of R1-Zero training via RL, they find that the increasing response length results not from self-reflection, but from RL optimizing well-designed rule-based rewards. Moreover, Yang et al. [904] demonstrate that the “aha moment” is externally marked by increased use of anthropomorphic language during self-reflection and a dynamic adjustment of uncertainty in response to problem difficulty. This process enables the model to maintain reasoning without succumbing to "Reasoning Collapse." Internally, it is characterized by a clear distinction between anthropomorphic traits and logical reasoning, with anthropomorphic language intensifying as the problem becomes more complex.